The electric field midway between two equal but opposite point charges is 736 N/C , and the distance between the charges is 15.0 cm . What is the magnitude of the charge on each?
Answer #1

use the equation for electric field to solve the problem

the equation for electric field due to charge is

E = kq/r^2

----------------------------------------------------------------------------------------------------

The net electric field midway between two equal but opposite point charges is

E net = kq/r^2 + kq/r^2

= 2kq/r^2

q = E net r^2/ 2k

= 736 (0.15 m/2)^2/ 2 *9 * 10^9

=2.3* 10^-10 C